### v3

# Cosmological constant as an integration constant

Abstract
The discrepancy between the observed value of the cosmological constant (CC) and its expected value from quantum field theoretical considerations motivates the search for a theory in which the CC is decoupled from the vacuum energy. In this article, we consider theories in which the CC is regarded as an integration constant. These theories include trace-free Einstein gravity, theories constructed from the Codazzi equation (which includes Cotton gravity and a gauge-gravity inspired theory), and conformal Killing gravity. We remark on a recent debate regarding the viability of Cotton gravity and find that while the Codazzi equation is indeed underdetermined, the solutions of the Codazzi equation trivialize to $\lambda g_{ab}$ on generic backgrounds, and that in principle, one can close the system with the divergence-free condition and an appropriate choice of initial data. We also propose a full variational principle (full in the sense that variations in all variables are considered) for each of the aforementioned theories that can incorporate the matter sector; in this manner, we can obtain the trace-free Einstein equations without a unimodular constraint.
Introduction — Since Einstein introduced the cosmological constant (CC) to his equations in 1917 [1], the debate about its nature has never stopped. Before the discovery of the accelerating expansion of the universe in 1999 [2, 3], the main consideration was why the CC is not exactly zero (See, for example, the famous review by Weinberg[4]). In the aftermath of the discovery of the accelerating expansion of the universe, the problem became more acute. The notion of dark energy introduced to account for this effect behaves like the CC. Indeed, in most analyses in cosmology, such as that of the cosmic microwave background, the CC is commonly invoked in the modeling. If true, then the observed value for the CC is tiny compared with the microscopic interpretation of the CC a la quantum field theory. Why is the CC not only nonzero but also so tiny? Such a seeming dilemma further motivates the reconstruction of Einstein gravity.
Such a reformulation should somehow decouple the CC from microscopic physics. One strategy is to seek a formulation of Einstein gravity in which the CC arises as a constant of integration. At the level of the field equations, there are several proposals in the literature. A proposal originating from Einstein [5] is that of the trace-free Einstein equations, which are obtained in unimodular gravity [6, *vanderBij:1981ym, *Buchmuller:1988wx, *Buchmuller:1988yn, *Henneaux:1989zc, 4] (see [11, *Alvarez:2023utn] for recent overviews of the topic), but can be considered on their own without the constraint on the metric determinant [13, 14]. Another proposal by Cook and Chen [15, 16], inspired by a set of third-order equations developed by Kilmister, Newman [17, *Kilmister:1961], and later by Yang [19] is to construct a gravity theory from the Bianchi identities and derivatives of the energy-momentum tensor. We focus primarily on the aforementioned approaches, though we mention here other proposals in which one may obtain the unconstrained Einstein field equations with the CC arising as an integration constant, such as the thermodynamical approach of Padmanabhan [20, 21] (see also [22]), a modified metric-affine theory [23, *Tapia:1998qt, *Tapia:1998as], and contributions from a three-form action  [^1] However, we note that in this case the vacuum energy still contributes to the cosmological term. [26, 27]. More recently, Harada proposed two theories of gravity, termed Cotton gravity [28] and the so-called conformal Killing gravity [29, 30]. Cotton gravity is constructed from the Cotton tensor, and it has been shown that one can rewrite the field equations of Cotton gravity as a Codazzi equation $\nabla_{[a}C_{b]c}=0$ [31, 32]; in this regard, Cotton gravity is in the same class as the earlier theory of Cook and Chen, as we will show that the field equations for the latter can also be reformulated as a Codazzi equation. Conformal Killing gravity, on the other hand, has a different structure altogether, as the field equations have the form of the defining equation for a gradient conformal Killing tensor.
Cotton gravity was recently the focus of a debate [33, 34, 35, 36] over its viability as a physical theory, as it was discovered that the symmetry-reduced equations of Cotton gravity are underdetermined. It is been argued in [34, 36] that while the equations of Cotton gravity are indeed underdetermined in cases of high symmetry (for instance, cosmological or spherically symmetric spacetimes), this is not necessarily the case in generic spacetimes. Similar concerns apply to the theory of Cook and Chen, as they share the same Codazzi structure. Another concern that applies to Codazzi type theories and conformal Killing gravity, is the fact that they are only defined at the level of the field equations; these theories suffer from the lack of a complete variational principle, as the variational principles proposed so far require holding some of the variables (such as the spacetime metric $g_{ab}$) fixed. One difficulty is that the energy-momentum tensor is obtained from variations of the metric tensor in the matter sector, but derivatives of the energy-momentum tensor appear in the field equations. The primary aim of this article is to shed light on the debate over the Codazzi equation and to propose a variational principle (including matter) for the trace-free Einstein equations, the Codazzi equation, and the conformal Killing equation without introducing additional constraints.
Trace-free Einstein equations — The trace-free Einstein equations may be obtained by assuming that the determinant of the metric is held fixed under the variation of the action; in this form, the theory is called unimodular gravity. However, the equations themselves can be written without assuming such a constraint [13, 14]:
$R_{\mu\nu}-\frac{1}{4}Rg_{\mu\nu}=\kappa(T_{\mu\nu}-\frac{1}{4}Tg_{\mu\nu}),\qquad\nabla^{\mu}T_{\mu\nu}=0,$ (1)
where $\underline{R}_{\mu\nu}$ is the Ricci tensor and $\nabla^{\mu}T_{\mu\nu}=0$ is imposed as an independent condition on the energy-momentum tensor. The contracted Bianchi identity $\nabla^{\mu}(R_{\mu\nu}-\tfrac{1}{2}Rg_{\mu\nu})=0$ combined with Eq.(1) implies $\nabla_{\mu}(R+\kappa T)=0$, which has the solution $R+\kappa T=4\bar{\Lambda}$, where $\bar{\Lambda}$ is an integration constant. Solving for $T$ and plugging this back into Eq. (1) yields:
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}+\bar{\Lambda}g_{\mu\nu}=\kappa T_{\mu\nu}.$ (2)
Later, we consider a variational principle that does not impose a unimodular constraint on the metric.
Bianchi gravity — The works [15, 16] consider a field theory in which the connection and curvature tensors are treated as analogues of the respective gauge field and field strength tensor of electrodynamics, yielding:
$\nabla_{\rho}R_{\mu\nu\sigma\tau}+\nabla_{\tau}R_{\mu\nu\rho\sigma}+\nabla_{\sigma}R_{\mu\nu\tau\rho}=0$ (3)
$\nabla_{\sigma}R^{\sigma\rho}{}_{\mu\nu}=\tfrac{\kappa}{2}J_{\mu\nu}{}^{\rho}$ (4)
where Eq. (3) is the Bianchi identity and $R_{\mu\nu\sigma\tau}$ is the curvature tensor. One postulates the source $J_{\mu\nu}{}^{\rho}$ to be integrable, so that for some symmetric rank-2 tensor $\bar{T}_{\mu\nu}$, it may be written as:
$J_{\mu\nu}{}^{\rho}=-q\nabla_{[\mu}\bar{T}_{\nu]}{{}^{\rho}}.$ (5)
The field equation (4) may be rewritten:
${\nabla}_{[\mu}R_{\nu]}{}{{}^{\rho}}-{\kappa}\nabla_{[\mu}\bar{T}_{\nu]}{{}^{\rho}}=0,$ (6)
where $\bar{T}_{\mu\nu}={T}_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}T$ is identified with the trace-reversal of the energy-momentum tensor. One class of solutions for Eq. (6) corresponds to the trace-reversed Einstein equation, up to some integration tensor $X_{\mu\nu}$:
$R_{\mu\nu}=\kappa({T}_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}T+{X}_{\mu\nu}),$ (7)
where $X_{\mu\nu}$ satisfies the following:
$X_{\mu\nu}=X_{\nu\mu},\qquad\nabla^{\mu}X_{\mu\nu}=0,\qquad\nabla_{[\mu}X_{\nu]\rho}=0.$ (8)
The tensor $X_{\mu\nu}$ is assumed to be determined by boundary and initial data; if the initial/boundary data yields $X_{\mu\nu}=\lambda g_{\mu\nu}$, then the CC $\lambda$ may be interpreted as an integration constant (we discuss later how such a solution arises from initial data).
One might imagine rewriting Eq. (6) as:
${\nabla}_{[\mu}K_{\nu]}{}{{}^{\rho}}=0,$ (9)
where:
$K_{\mu\nu}=R_{\mu\nu}-\kappa({T}_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}T+{X}_{\mu\nu}).$ (10)
Equation (9) has the form of a Codazzi equation [31, 32], which is the subject of recent interest in the Cotton gravity program [28].
Cotton gravity — Cotton gravity is based on the Cotton tensor:
$C_{abc}:=\nabla_{b}R_{ac}-\nabla_{c}R_{ab}-\tfrac{1}{6}\left(g_{ac}\nabla_{b}R-g_{ab}\nabla_{c}R\right)$ (11)
Defining a new tensor $\mathscr{T}_{abc}$
$\mathscr{T}_{abc}:=\nabla_{b}T_{ac}-\nabla_{c}T_{ab}-\tfrac{1}{6}\left(g_{ac}\nabla_{b}T-g_{ab}\nabla_{c}T\right),$ (12)
one can write the governing equation for Cotton gravity as:
$C_{abc}=2\kappa\mathscr{T}_{abc}.$ (13)
This equation can be rewritten in the form:
${\nabla}_{[a}\mathcal{C}_{b]}{}{{}^{c}}=0,$ (14)
where:
$\mathcal{C}_{ab}=G_{ab}-\kappa T_{ab}-\frac{1}{3}(G-\kappa T)g_{ab}+X_{ab}.$ (15)
“Conformal” Killing gravity — The equations for the so-called conformal Killing gravity theory proposed in [29] may be written in the form [30]:
$6(K_{bc;a}+K_{ca;b}+K_{ab;c})=g_{bc}K_{,a}+g_{ca}K_{,b}+g_{ab}K_{,c},$ (16)
$K_{ab}:=\kappa T_{ab}-G_{ab}.$
The equation above has the form of the equation for a conformal Killing tensor:
$K_{bc;a}+K_{ca;b}+K_{ab;c}=u_{a}g_{bc}+u_{b}g_{ca}+u_{c}g_{ab},$ (17)
where $u_{a}$ is a one-form; $K_{ab}$ is a symmetric Killing tensor when $u_{a}=0$. However, as pointed out in [37] a simple redefinition $\bar{K}_{ab}:=6K_{ab}-g_{ab}K$ reveals that any nontrivial solution of Eq. (16) requires the existence of a Killing tensor given by $\bar{K}_{ab}$. It follows that the solutions of Eq. (16) either trivialize to $K_{ab}=-\lambda g_{ab}$, or admit a Killing tensor. For this reason, it is perhaps more appropriate to refer to this theory as “Killing tensor gravity”, but we retain here the common terminology in the literature. Moreover, we note that perturbations of Killing tensor solutions satisfying $K_{ab}\neq-\lambda g_{ab}$ must also admit a Killing tensor; since our universe (being inhomogeneous and anisotropic on small scales) does not have the required symmetries to admit a global Killing tensor, we regard such solutions as unphysical.
The Codazzi equation — Equations (9) and (14) have the form of a Codazzi equation:
${\nabla}_{[a}C_{b]}{}{{}^{c}}=0,$ (18)
where $C_{ab}=C_{ba}$ is termed a Codazzi tensor, following the terminology of [31, 32]. One can perhaps imagine a class of theories defined by an equation of the form in Eq. (18). Now there is some debate over whether equations of this type are predictive [33, 34, 35, 36]; in particular, it has been claimed in [35] that the Codazzi equations are underdetermined, and in a responding comment it was claimed that this only occurs for highly symmetric spacetimes [36]. In the following, we attempt to shed some light on this dispute.
Contracting Eq. (18), one obtains:
${\nabla}_{c}C_{a}{}{{}^{c}}-{\nabla}_{a}C=0.$ (19)
where $C:=C_{a}{}{{}^{a}}$. Taking the divergence of Eq. (18), one obtains a wave-like equation:
$\Box C_{a}{}{{}^{b}}-\nabla_{a}{\nabla}_{c}C{}^{cb}-R_{ac}C^{cb}-R_{acd}{{}^{b}}C^{cd}=0.$ (20)
and the divergence of Eq. (19) is:
$\Box C-{\nabla}_{a}{\nabla}_{b}C^{ab}=0.$ (21)
Of course, Eq. (20) is underdetermined due to the term $\nabla_{a}{\nabla}_{c}C{}^{cb}$. In an orthonormal basis, the derivative $\partial^{2}_{t}C{{}_{t}}^{b}$ disappears from Eq. (20), so that the time evolution of the components $C{{}_{t}}^{b}$ is ill-defined—the equations are indeed underdetermined. However, Eq. (20) reveals a straightforward resolution; one should supply the Codazzi equation with a constraint equation, and a natural one is the divergence-free condition $\nabla_{a}C^{ab}=0$. Of course, an attempt to directly solve the Codazzi equations supplemented with the $\nabla_{a}C^{ab}=0$ may still lead to ambiguities, but such ambiguities correspond to freedom in choosing initial data for $C_{ab}$ and $\partial_{t}C_{ab}$ for the equivalent system:
$\Box C_{a}{}{{}^{b}}-R_{ac}C^{cb}-R_{acd}{{}^{b}}C^{cd}=0,\qquad\nabla_{a}C^{ab}=0.$ (22)
Now Eq. (20) follows from Eq. (18). It is perhaps appropriate to consider the application of the D’Alembertian to the following quantity:
$Q_{ab}{{}^{c}}:=2{\nabla}_{[a}C_{b]}{}{{}^{c}},$ (23)
which after using Eq. (20) has the form:
$\Box Q_{cde}=$ $R_{cdeb}\nabla_{a}C^{ab}+2(\nabla^{b}C^{a}{}_{[c})R_{d]bea}+4\nabla_{[e}R_{a][c}C_{d]}{}^{a}$ (24)
$+2C^{ab}\nabla_{[d}R_{c]aeb}-Q_{d}{}^{a}{}_{e}R_{ca}-Q_{ace}R_{d}{}^{a}$
$-2Q^{ab}{}_{e}R_{cadb}+Q_{d}{}^{ab}R_{caeb}-Q_{c}{}^{ab}R_{daeb}.$
Note that the right-hand side contains at most first derivatives of $C_{ab}$ (and also that the divergence-free condition removes the first term). One can then regard Eq. (24) as a wave equation for $Q_{cde}$, with $C_{ab}$ being an independent variable satisfying Eq. (20) and the constraint (24). Now consider initial data consistent with Eq. (18), so that $Q_{cde}=0$ (but not its derivatives). For nonvanishing curvature, the right-hand side will generically produce a source unless $C_{ab}=\lambda g_{ab}$ on a Cauchy hypersurface and $\nabla_{c}C_{ab}=0$. This indicates that if $C_{ab}\neq\lambda g_{ab}$ and $\nabla_{c}C_{ab}\neq 0$ satisfies Eq. (24) on a Cauchy surface, the time evolution of $C_{ab}$ under Eq. (20) will in general generate violations of Eq. (24). There may also be special cases (highly symmetric spacetimes, for instance) in which Eq. (18) may admit nontrivial solutions $C_{ab}\neq\lambda g_{ab}$, but one cannot expect the existence of solutions to Eq. (18) for generic initial data. It should be noted that Eq. (24) and the conclusions outlined in the previous paragraph are independent of the constraint $\nabla_{a}C^{ab}=0$.
In short, Eq. (20) indicates that the quantity $Q_{ab}{{}^{c}}$ evolves according to Eq. (24), and that generically, the condition $Q_{ab}{{}^{c}}=0$ can only be maintained for initial data also satisfying $C_{ab}=\lambda g_{ab}$ and $\nabla_{c}C_{ab}=0$, where $\lambda$ is an arbitrary constant; other solutions will likely require background geometries with special properties, as argued in [34, 36]. We find that both sets of authors, [33, 35] and [34, 36], present valid points: theories based on the Codazzi tensors are indeed underdetermined, but the space of solutions is nonetheless highly restricted in generic settings.
Variational principle — The variational principles presented in the literature for the trace-free Einstein equations [4] and theories admitting a Codazzi equation formulation (Bianchi gravity [15, 16] and Cotton gravity [28]) typically require holding certain quantities fixed; for the trace-free Einstein equations, the metric determinant is held fixed (corresponding to an additional constraint), and for the Codazzi equation, the metric is held fixed (as variations in the metric will generally introduce additional unwanted constraints). We therefore seek a general variational principle that yields the desired field equations (including contributions from the matter sector) without unwanted constraints under a full variation—that is, a variation with respect to all variables appearing in the action.
Consider a simple action of the form for the case of the Codazzi equation:
$S_{C}:=$ $\int_{\Omega}\sqrt{|g|}d^{4}x[Y^{abc}\nabla_{[a}C_{b]c}+\Psi_{abcdef}Y^{abc}Y^{def}].$ (25)
[…]
$\nabla_{[a}C_{b]c}+2\Psi_{abcdef}Y^{def}=0,$ (26)
which reduces to the Codazzi equation when $Y^{abc}=0$. The variation of the action $S_{C}$ with respect to $g^{ab}$ and $Z_{ab}$ is rather nontrivial, but ultimately consists of terms proportional to products of $Y^{abc}$ and its derivatives; variations with respect to $g^{ab}$ and $Z_{ab}$ vanish when $Y^{abc}=0$. To independently enforce the divergence-free constraint, one may add to the action the following:
$S_{D}:=$ $\int_{\Omega}\sqrt{|g|}d^{4}x[U^{b}\nabla^{a}C_{ab}+\psi_{ab}U^{a}U^{b}],$ (27)
where $U^{a}$ is an auxiliary field and $\psi_{ab}=\psi_{ba}$ is a Lagrange multiplier. The variation of $S_{D}$ with respect to $\psi_{ab}$ implies $U^{a}=0$, and the variation with respect to $U^{a}$ yields:
$\nabla^{a}C_{ab}+2\psi_{ab}U^{a}=0.$ (28)
When $U^{a}=0$, the above reduces to the divergence-free condition, and the variation of $S_{D}$ with respect to the remaining variables vanishes.
It is straightforward to apply this strategy to the other theories considered in this article. For conformal Killing gravity, we construct the action:
$S_{K}:=$ $\int_{\Omega}\sqrt{|g|}d^{4}x[Y^{abc}\nabla_{(a}\bar{K}_{bc)}+\Psi_{abcdef}Y^{abc}Y^{def}].$ (29)
where $\bar{K}_{ab}:=6K_{ab}-g_{ab}K$. For the trace-free Einstein equations, we have:
$S_{TF}:=$ $\int_{\Omega}\sqrt{|g|}d^{4}x[Y^{ab}\left(\underline{R}_{ab}-\kappa\underline{Z}_{ab}\right)+\Psi_{abcd}Y^{ab}Y^{cd}],$ (30)
where $\underline{R}_{ab}:=R_{ab}-\tfrac{1}{4}Rg_{ab}$ and $\underline{Z}_{ab}:=R_{ab}-\tfrac{1}{4}Zg_{ab}$. We emphasize that no constraint on the metric determinant is required.
The matter sector is a somewhat delicate matter. Naively, one can add a term to the matter action:
$S^{\prime}_{M}=S_{M}+\int_{\Omega}\sqrt{|g|}d^{4}x(\tilde{Z}_{ab}g^{ab}),$ (31)
so that the variation with respect to $g^{ab}$ yields $\tilde{Z}_{ab}=T_{ab}-\tfrac{1}{2}Tg_{ab}$. However, the difficulty with such a strategy is that the variation $\delta\tilde{Z}_{ab}$ will place a nontrivial constraint on $g^{ab}$ unless it becomes a surface term $\delta\tilde{Z}_{ab}=\nabla_{c}\phi^{c}_{ab}$, or otherwise trivializes. However, the only known tensor with this property is a Ricci tensor formed from a metric-compatible connection, in which case we recover (by way of a Palatini variation) the Einstein equations with a specified CC, or the corresponding Einstein-Cartan equations if fermions are included.
Consider the action:
$S_{\bar{M}}=\int_{\Omega}\sqrt{|g|}d^{4}x\biggl{[}(g^{ab}-h^{ab})Z_{ab}+(L_{h}-L_{g})\mathcal{E}+\Xi+J\cdot\zeta\cdot J\biggr{]},$ (32)
where $\phi$ is a matter field (assumed to be nonspinorial and minimally coupled), $\chi$ is an extra field that is equal to $\phi$ on shell, $h^{ab}$ and $J$ are auxiliary fields, $\zeta$ is a Lagrange multiplier [^3] In particular, one assumes $\zeta$ to be defined so that $J\cdot\zeta\cdot J$ is a scalar., and the following have been defined:
$L_{h}:=L(\chi,\partial_{a}\chi,h^{ab}),\qquad L_{g}:=L(\varphi,\partial_{a}\varphi,g^{ab}),$ (33)
$\Xi:=(\varphi-\chi)\cdot J,\qquad\mathcal{E}(g^{ab},\underline{h}_{ab}):=\exp\left[\sqrt{g^{ab}\underline{h}_{ab}}-2\right].$
The variation with respect to $Z_{ab}$ yields $h^{ab}=g^{ab}$, the variation with respect to $\zeta$ yields $J=0$, and the variation with respect to $J$ yields $\varphi=\chi$ when evaluated on $J=0$. The variation of $\mathcal{E}$ has the form:
$\delta\mathcal{E}=\frac{\exp\left\{\sqrt{g^{ab}\underline{h}_{ab}}-2\right\}}{\sqrt{g^{ab}\underline{h}_{ab}}}\biggl{[}\delta g^{ab}\underline{h}_{ab}-g^{cd}\underline{h}_{ac}\underline{h}_{bd}\delta h^{ab}\biggr{]},$ (34)
which on $h^{ab}=g^{ab}$ simplifies to:
$\delta\mathcal{E}=\frac{1}{2}g_{ab}\left(\delta g^{ab}-\delta h^{ab}\right),$ (35)
Similarly, the Lagrangian variations satisfy:
$\delta L_{h}$ $=\frac{\partial L_{h}}{\partial h^{ab}}\delta h^{ab}+\frac{\partial L_{h}}{\partial\chi}\delta\chi+\frac{\partial L_{h}}{\partial(\partial_{a}\chi)}\delta\partial_{a}\chi,$ (36)
$\delta L_{g}$ $=\frac{\partial L_{g}}{\partial g^{ab}}\delta g^{ab}+\frac{\partial L_{g}}{\partial\varphi}\delta\varphi+\frac{\partial L_{g}}{\partial(\partial_{a}\varphi)}\delta\partial_{a}\varphi.$
One finds that the variations with respect to $g^{ab}$ and $h^{ab}$ yield two constraints that coincide when evaluating on $g^{ab}=h^{ab}$ and $\varphi=\chi$:
$Z_{ab}=\frac{\partial L_{h}}{\partial h^{ab}}-\frac{1}{2}h_{ab}L_{h}=\frac{\partial L_{g}}{\partial g^{ab}}-\frac{1}{2}g_{ab}L_{g},$ (37)
which makes use of the fact that for $g^{ab}=h^{ab}$ and $\varphi=\chi$, the integrand of Eq. (32) is constructed to vanish. The variations with respect to $\varphi$ and $\chi$ yield matter field equations that coincide on $g^{ab}=h^{ab}$ and $\varphi=\chi$, and reduce to the usual field equations obtained from $L_{g}$ after applying the constraint $J=0$.
Discussion — In this article, we considered some theories in which the CC arises as an integration constant, including trace-free Einstein gravity, the Codazzi equation, and conformal Killing gravity. We revisited a recent debate [33, 34, 35, 36] concerning the issue of nonuniqueness of solutions to equations of Cotton gravity in symmetry-reduced settiongs. By taking additional derivatives of the Codazzi equations, we find that solutions of the Codazzi equations must satisfy at least one of two conditions: either the solutions coincide with that of the Einstein equations (with arbitrary CC), or that certain contractions of the curvature tensor, the Codazzi tensor, and their derivatives must vanish (in particular, all terms on the right-hand side of Eq. (24) must vanish). The latter can be satisfied in situations of high symmetry, even if the solutions do not satisfy the Einstein equations, but if we regard such non-Einstein solutions as physical, the right-hand side of Eq. (24) must vanish exactly even under perturbations. It is unlikely that such conditions can be satisfied unless the perturbations are highly restricted, and one may in this manner argue that such non-Einstein solutions are unphysical. We find that the Codazzi equations can in principle be closed with the addition of the divergence-free constraint so that such non-Einstein solutions can be excluded at the level of initial data.
We formulated a rather general variational principle that can yield the trace-free Einstein equations, the Codazzi equation, and the equations of conformal Killing gravity. An action for trace-free gravity without a unimodular constraint [38] was recently proposed (another possible approach is described in [23, *Tapia:1998qt, *Tapia:1998as]), but the matter sector was not explicitly incorporated in their treatment; our action (30) provides an alternative that can readily incorporate the matter sector via the action (32). The appearance of Lagrange multipliers and auxiliary fields in the actions corresponds to introducing additional nondynamical degrees of freedom—one might prefer an action that depends exclusively on dynamical degrees of freedom. Moreover, there is some question of whether auxiliary fields in gravity theories are pathological [39, 40, 41]. Regarding the latter, there may be instances in which some of the pathologies (particularly those associated with sharp density gradients) can be mitigated [42, 43, 44], and one can perhaps regard auxiliary fields as effective descriptions for underlying dynamical degrees of freedom (with large masses). Regarding the former concern, it is well-known that one can express the Lagrangian for $f(R)$ theories in the O’Hanlon [45] form $\psi R-V(\psi)$ (see also [46, 47, 48, 49, 50] and references therein), where $\psi$ is an auxiliary field, and one can easily recover the $f(R)$ Lagrangian from the O’Hanlon Lagrangian by algebraic elimination. Of course, a naive algebraic elimination would trivialize the actions considered here, but the point is that auxiliary fields can provide a starting point for constructing an action exclusively from dynamical degrees of freedom. To clarify, we are not suggesting that the variational principle we have proposed is fundamental—indeed, one can construct a variational principle for virtually any set of consistent field equations in this manner. [^4] Presumably one could attempt to use the same strategy to construct a variational principle for inconsistent field equations, but this will result in an action that does not admit extrema. Instead, we argue that it can serve as a first step towards the construction of a more fundamental action for these theories.
Ghosts are a concern; for instance, the Codazzi equation contains up to three derivatives of the metric. One might also recognize that the duplicate matter degrees of freedom $\chi$ in Eq. (32) are necessarily ghosts. At the classical level, there is no issue in the matter sector, as the matter field equations ultimately reduce to those identical to that obtained in the standard variational principle—there is no runaway behavior (see also [51]). In the gravitational sector, one can place restrictions on the initial data such that the system becomes dynamically equivalent to the Einstein equations with a CC. Regarding the Hamiltonian, we note that on shell, both the matter and gravitational actions vanish in our approach, and it is not too difficult to convince oneself that the total matter Hamiltonian must also vanish on shell. Though there is no issue at the classical level, the appearance of ghosts may be of particular concern for the quantization of the theory. Quantization is beyond the scope of the present article (see the discussion in [16] regarding ghosts in higher curvature theories), but we leave the reader with a couple of remarks. First, as argued in the preceding paragraph, the action presented in this article may require further modification at high energies, where the structure of the resulting action may differ significantly. Second, the vanishing of the Hamiltonian indicates that these theories suffer from the problem of time, in which the Hamiltonian operator does not yield the time evolution of the quantum state of the field, a well-known problem in canonical approaches to quantum gravity [52, *Isham:1992ms].
Acknowledgments — We thank Marco de Cesare, Naresh Dadhich, and Victor Tapia for helpful comments and references. PC and JCF acknowledge support from the Taiwan National Science and Technology Council (NSTC) under project numbers 110-2112-M-002-031 and 112-2811-M-002-132, and also from the Leung Center for Cosmology and Particle Astrophysics (LeCosPA), National Taiwan University.